Reaction Stoichiometry os Over the years, the thermite reaction has been used for weld- ing railroad rails, in incendiary bombs, and to ignite solid-fuel rocket motors, The reaction is Fed,() + 2Al(s)-→ 2Fe() + A120,G) What masses of iron I) oxide and aluminum must be used to produce 15.0 g iron? What is the maximum mass of aluminum oxide that could be produced? 106 The reaction between potassium chlorate and red phosphorus takes place when you strike a match on a matchbox. If you were to react 52.9 g of potassium chlorate (KCIOs) with ex- cess red phosphorus, what mass of tetraphosphorus decaoxide P,O10) could be produced? KCIO3(s) + P4(s)-→ P,0。(s) + KCl(s) (unbalanced)
Answer #1

105)

a)

mass of Fe = 15 g

molar mass of Fe = 55.85 g/mol

mol of Fe = (mass)/(molar mass)

= 15/55.85

= 0.268577 mol

According to balanced equation

mol of Fe2O3 required = (1/2)* moles of Fe

= (1/2)*0.268577

= 0.134288 mol

mass of Fe2O3 = number of mol * molar mass

= 0.134288*215.55

= 28.9 g

Answer: 28.9 g

b)

According to balanced equation

mol of Al required = moles of Fe

= 0.268577 mol

mass of Al = number of mol * molar mass

= 0.268577*26.98

= 7.25 g

Answer: 7.25 g

c)

According to balanced equation

mol of Al2O3 formed = (1/2)* moles of Fe

= (1/2)*0.268577

= 0.134288 mol

mass of Al2O3 = number of mol * molar mass

= 0.134288*128.94

= 17.3 g

Answer: 17.3 g

I am allowed to answer only 1 question at a time